# The Atlas of In-Context Learning: How Attention Heads Shape In-Context Retrieval Augmentation

### v1

@@ 1 Introduction
Viewing a prompt as a composition of informational components, we show that on question answering tasks, certain attention heads perform various operations on the prompt at different stages of inference and layers, as depicted in Figure 1. Our method identifies two groups of heads based on their functions: parametric heads that encode relational knowledge [50, 22] and in-context heads responsible for processing information in the prompt. Further, as in-context heads need to understand which prompt components to process and how, we hypothesize that they specialize to fill their respective roles.
@@
Building on these insights, we probe for sources of knowledge used during retrieval-augmented generation (RAG) and show where it is localized within the input context. Our attempt shows promising results, serving as an initial step towards more transparent RAG systems.
Overall, our contributions can be summarized as follows:
  • We describe an attribution-based method to determine attention heads that play a key role during in-context retrieval augmentation, revealing that they operate on the prompt in a distinct manner. Our method can thus be used to reliably categorize attention heads into in-context and parametric heads.
@@ 2 Related Work
Retrieval Augmentation    Retrieval-augmented language models (RALMs) [34, 38] address inherent limitations of LMs by providing external knowledge sources. Despite this advantage, issues such as discrepancies between contextual and parametric knowledge may occur [41, 74]. Some works have studied mechanisms behind knowledge preference in RALMs [77, 31, 48, 44], but they mostly focus on simple domains where the answer is limited to a single word. Besides, RALMs cannot guarantee faithful answer attribution [^1] Here, the term answer attribution means the use of external documents to support the generated response, which is different from feature attribution used throughout this work to describe interpretability techniques. to contextual passages [20], which necessitates a shift to interpretability.
@@
In-Context Learning    Numerous works have studied ICL since its introduction. Liu et al. [40] studied what constitutes a good example for demonstrations. Dai et al. [16] suggested that ICL can be understood as an implicit fine-tuning. ICL is a general phenomenon, although it is commonly assumed to be unique to autoregressive models [54]. At the component level, ICL is primarily associated with induction heads [17, 47]. However, recent findings showed that certain heads can also be compressed to form FVs that represent a demonstrated task [26, 62]. Yin and Steinhardt [76] investigated the connection between these heads and induction heads, showing that they are distinct from each other and how models’ ICL performance is mainly driven by the former. ICL can also be viewed as a combination of meta-learning and retrieval [45]. In that regard, we study the latter perspective to understand its mechanism, with a focus on the retrieval augmentation paradigm.
@@ 3 Background and Preliminaries
An alternative would be to use feature attribution methods [9], as they trace the contribution of each input element to the model’s output prediction. Propagation-based feature attribution [58, 8, 59] especially takes the entire computation path into account, which can be used to characterize attention heads [64] or identify latent concepts [1]. Furthermore, feature attribution is able to estimate causal relations [21] and thus enables to observe how a specific attention head affects a model’s prediction.
@@ 3.2 Attention Mechanism
While the original formulation of the multi-head attention mechanism [63] concisely summarizes the parallel computation of attention heads, our goal is to isolate their individual contributions. To this end, we reformulate the equations to make the influence of each head more explicit [17, 18]. Let $\mathbf{X}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{S})\in\mathbb{R}^{d\times S}$ denote the matrix of hidden token representations for a sequence $S$ with dimension $d$, and suppose our model employs $H$ parallel heads, each of dimension $d_{h}=d/H$. Then, the computation of the multi-head attention layer can be reformulated into $H$ complementary operations, where each head $h$ produces an intermediate attention output $\mathbf{z}_{i}^{h}\in\mathbb{R}^{d_{h}}$:
@@
where $\mathbf{A}^{h}_{i,j}$ is the attention weight of token $i$ attending to token $j$, and $\mathbf{W}_{V}^{h}\in\mathbb{R}^{d_{h}\times d}$ is the per-head value projection. The final output is obtained by multiplying the intermediate output of each head with their corresponding output projection matrix $\mathbf{W}_{O}^{h}\in\mathbb{R}^{d\times d_{h}}$, followed by summing:
@@ 4 Experimental Setup
Similar to Petroni et al. [51], we consider different types of contextual information to see how they affect in-context and parametric heads. We use oracle contexts as they are always relevant to the question and contain the true answer. In addition, we use counterfactual contexts as they contain information that is usually not seen during pre-training and fine-tuning stages, thus forcing models to rely on the text to answer the question correctly. Oracle context is often not available; therefore we also use Dense Passage Retriever (DPR) [33] with a Wikipedia dump from December 2018 as our retrieval corpus. For simplicity, we only select the top one retrieved document. We show results for oracle and counterfactual contexts in the main paper and retrieved context in Appendix B.
@@ 5 Localization of In-Context and Parametric Heads
We examine whether in-context and parametric heads are independent of each other. As expected, ablating parametric heads has little influence to the model’s performance in our open-book setting. Interestingly, this leads to a slight performance increase on NQ-Swap with counterfactual contexts, which suggest that the ablation forces the model to rely more on the given context instead of its own parametric knowledge. Surprisingly, ablating in-context heads in the closed-book setting incurs a non-negligible performance reduction. This is likely due to the influence in-context heads have when processing the input prompt. We explore this in the next section.
@@ 6 Functional Roles of In-Context and Parametric Heads
To answer correctly, the model must map the intensional frame onto the knowledge triple $(s,r,o^{*})$, where $s$ is the subject (“Mike Tyson”), $r$ is the semantic relation (the predicate specified by the question, here “has worked as”), and $o^{*}$ is the object (the yet to be determined answer, “boxer” or “firefighter”). Depending on where the answer resides, $o^{*}$ may be retrieved from the model’s parametric memory (${\color[rgb]{0.8515625,0.265625,0.359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.265625,0.359375}o^{p}}$) or from the context (${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}o^{c}}$). By treating $(s,r,o^{*})$ as the complete task specification, we analyze how in-context and parametric heads specialize both to comprehend the intensional frame and to retrieve the object $o^{*}$ needed to generate the correct answer.
@@ 6.1 Disentangling the Functionality of In-Context Heads
Finally, to obtain the sets of specialized task and retrieval heads, we rank heads by their aggregated relevance and select the top $K$, a hyperparameter that will be determined in the next section.
$\mathcal{H}_{\text{task}}=\bigl{\{}\mathrm{argsort}(\rho^{h}_{\text{task}})_{\mathrm{desc}}\bigr{\}}_{n=1}^{K},\quad\mathcal{H}_{\text{ret}}=\bigl{\{}\mathrm{argsort}(\rho^{h}_{\text{ret}})_{\mathrm{desc}}\bigr{\}}_{n=1}^{K}.$ (9)
Results   We compute the task relevance score $\rho^{h}_{\text{task}}$ and the retrieval relevance score $\rho^{h}_{\text{ret}}$ over NQ-Swap with counterfactual contexts to minimize influences of parametric heads, and aggregate their distributions across the model layers. In Figure 1, we observe that $\rho^{h}_{\text{task}}$ initially increases in the early layers where few parametric heads are located, suggesting that early parametric heads enrich the question with relational knowledge.
The relevance peaks in the middle layers, where in-context heads dominate, aligning with the transition to a more context-dependent reasoning. In contrast, the retrieval relevance score $\rho^{h}_{\text{ret}}$ peaks in deeper layers, reflecting the point where the model extracts the final answer object ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}o^{c}}$. Figure 3 further illustrates the sorted average difference $\mathcal{D}$ between open-book and closed-book settings for all heads, alongside the top 40 task heads $\mathcal{H}_{\text{task}}$ and retrieval heads $\mathcal{H}_{\text{ret}}$. We observe that retrieval heads and task heads consistently rank among the highest-scoring in-context heads, emphasizing the critical role of in-context heads for retrieval augmented generation.
@@ 6.2 Causal Effects of In-Context and Parametric Heads
An important question is whether the heads we identify truly reflect their assigned functionalities. We examine this by investigating their causal effects on the answer generation process.
Experiments   We conjecture that task heads $\mathcal{H}_{\text{task}}$ encode the intensional frame $(s,r,o^{*})$ and that parametric heads $\mathcal{H}_{\text{param}}$ contain information of the subject $s$, which depending on the training data may or may not include ${\color[rgb]{0.8515625,0.265625,0.359375}\definecolor[named]{pgfstrokecolor}{rgb}{0.8515625,0.265625,0.359375}o^{p}}$. On the other hand, retrieval heads $\mathcal{H}_{\text{ret}}$ search for ${\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}o^{c}}$, allowing them to copy any tokens from the context verbatim, without being restricted to only plausible answers. For task and parametric heads, we compute FVs [^3] A function vector can be defined as a sum of averaged heads outputs over a task [62] or computed individually [26]. Here, we consider a function vector to be an individual output of a task or parametric head. for each head and insert them into various settings to trigger the execution of their functions. Following Wu et al. [71], we opt for a needle-in-a-haystack (NIAH) setting for retrieval heads and determine their ability to retrieve relevant information from the context by modifying the attention weights. We conduct our analysis on the biography dataset described in §4 and measure the recall score [3]. For comparison, we also consider random heads for FV extraction and attention modification. See Appendix C for additional details and results.
@@
Models $\mathcal{H}_{\text{task}}$ $\mathcal{H}_{\text{param}}$ $\mathcal{H}_{\text{ret}}$
@@ 7 Source Tracking
Our experiment in §6.2 demonstrates that retrieval heads reliably perform verbatim copying of text spans when their corresponding attention maps focus on the retrieved tokens. We now aim to investigate if we can (i) detect when retrieval heads initiate the copying process for the first answer token (i.e., whether a token is derived from external contexts rather than from the model parameters), and (ii) accurately localize its position within that context using the attention maps. To this end, we train a linear probe on NQ-Swap with counterfactual contexts. Each retrieval head’s output at the last token’s position $\mathbf{z}_{S}^{h}$ is decoded via logit lens [46], converting each head’s activation into a score for token $t\in\mathbb{N}$ using the model’s unembedding matrix $\mathbf{W}_{U}\in\mathbb{R}^{|V|\times d}$ and layer normalization $\mathrm{LN}(\cdot)$:
@@
Next, we train the probe by learning weights $\{w_{h}\}_{h\in\mathcal{H}_{\mathrm{ret}}}$. For the localization task, we aggregate each head’s attention map $A_{S,j}^{h}$ using the probe’s weight and the logit lens score, as their attention maps pinpoint the retrieved token. The source token position is predicted as the index with the highest aggregated attention weight. For details, please refer to Appendix D. Additionally, we use a standard AttnLRP backward pass from the model output to compute an input heatmap as a baseline for comparison.
@@
Models ROC AUC Attention Localization AttnLRP Localization
@@ 8 Conclusion
We propose a method to explore the inner workings of ICL for retrieval-augmentation, revealing in-context and parametric heads that operate on the input prompt in a distinct manner and find that in-context heads can specialize to read instructions or retrieve relevant information. We study the roles of the identified heads by converting them into FVs or modifying their weights, showing how they can affect the generation process. Finally, we present a probe to precisely and efficiently track for knowledge provenance, opening up a path towards more interpretable retrieval-augmented LMs.
Limitations    In this work, we focus our investigation on attention heads since they are primarily associated with ICL. However, how they interact with components in MLP modules e.g., knowledge neurons [15] to induce functions remains an open question. We leave this avenue for future work.
@@ Appendix A Appendix
Llama-3.1-8B-Instruct is released under the Llama 3.1 Community License. Gemma-2-9B-it is released under the Gemma license agreement. Mistral-7B-Instruct-v0.3 is released under Apache 2.0. NQ-Swap and TriviaQA that we use are derived from MRQA [19], which is released under the MIT license. We construct the biography dataset using Wikidata Query Service, [^4] https://query.wikidata.org/ which is available under CC0.
Datasets
For NQ-Swap, we use the preprocessed data and split available on HuggingFace [^5] https://huggingface.co/datasets/pminervini/NQ-Swap (4,746 examples). As for TriviaQA, we use the dev split from the MRQA repository [^6] https://github.com/mrqa/MRQA-Shared-Task-2019 (7,785 examples).
@@
ORIG Vladimir Vapnik was educated in V.A. Trapeznikov Institute of Control Sciences. Vladimir Vapnik was born 06 December 1936. Vladimir Vapnik worked as computer scientist. Vladimir Vapnik was born in Tashkent.
PARA Vladimir Vapnik was born on 06 December 1936 in Tashkent, a city that would later shape his life’s work. As a young man, Vapnik was fascinated by the potential of machines to learn and adapt. He went on to study at the V.A. Trapeznikov Institute of Control Sciences, where he was exposed to the latest advancements in computer science and artificial intelligence. It was here that Vapnik’s passion for machine learning truly took hold. After completing his studies, Vapnik went on to become a renowned computer scientist, making groundbreaking contributions to the field. His work on support vector machines and the Vapnik-Chervonenkis dimension would have a lasting impact on the development of machine learning algorithms. Throughout his career, Vapnik has been recognized for his innovative thinking and dedication to advancing the field of computer science. His legacy continues to inspire new generations of researchers and scientists.
@@
All the experiments were conducted on 2 x 24 GB RTX4090 and 4 x 40 GB A100. Computing difference score $\mathcal{D}$ to identify in-context and parametric heads takes about 8 hours. Ablating both heads on NQ-Swap and TQA takes about 3 hours for each run. Patching task and parametric FVs takes about 12 hours on average for each model, while the NIAH experiment with retrieval heads takes about 8 hours. Our source tracking experiments consume about 4 hours.

### v2

@@ 1 Introduction
On question answering tasks, we show that by viewing a prompt as a composition of informational components, certain attention heads perform various operations on the prompt at different stages of inference and layers (Figure 1). Our method identifies two groups of heads based on their functions: parametric heads that encode relational knowledge [57, 27] and in-context heads responsible for processing information in the prompt. Further, as in-context heads need to understand which prompt components to process and how, we hypothesize that they specialize to fill their respective roles.
@@
Building on these insights, we probe for sources of knowledge used during retrieval-augmented question answering and show where it is localized within the input context. Our attempt shows promising results, serving as an initial step towards more transparent retrieval-augmented generation (RAG) systems. Overall, our contributions can be summarized as follows:
  • We describe an attribution-based method to determine attention heads that play a key role during in-context retrieval augmentation for question answering, revealing that they operate on the prompt in a distinct manner. Our method can thus be used to reliably categorize attention heads into in-context and parametric heads.
@@ 2 Related Work
Retrieval Augmentation    Retrieval-augmented language models (RALMs) [40, 45] address inherent limitations of LMs by providing external knowledge sources. Despite this advantage, issues such as discrepancies between contextual and parametric knowledge may occur [48, 87]. Some works have studied mechanisms behind knowledge preference in RALMs [91, 55, 51], but they focus on simple domains. On the other hand, those that explored more complex domains mostly only analyzed RALMs’ behavior at the output level [14, 85, 71]. Closely related to our work is that of Jin et al. [37], where the authors also discovered in-context and parametric heads. Compared to them, we show that interactions among in-context heads are subtly more complex, since they can specialize into task or retrieval heads. Our approach also allows to demonstrate their roles in shaping the model’s representation along with parametric heads, by transforming them into FVs or modifying their respective weights.
Another disadvantage of RALMs is that they cannot guarantee faithful answer attribution [^1] Here, the term answer attribution means the use of external documents to support the generated response, which is different from feature attribution used throughout this work to describe interpretability techniques. to contextual passages [25], which necessitates a shift to interpretability. Recent efforts to address this include leveraging contrastive gradient attribution [60, 89], fitting surrogate models [16], and treating attention weights as features [17]. In relation to this, we train a probe based on retrieval heads to track for knowledge provenance.
@@
In-Context Learning    Numerous works have studied ICL since its introduction. Liu et al. [47] studied what constitutes a good example for demonstrations. Dai et al. [20] suggested that ICL can be understood as an implicit fine-tuning. ICL is a general phenomenon, although it is commonly assumed to be unique to autoregressive models [63]. At the component level, ICL is primarily associated with induction heads [21, 54]. However, recent findings showed that certain heads can also be compressed to form FVs that represent a demonstrated task [31, 74]. Yin and Steinhardt [90] investigated the connection between these heads and induction heads, showing that they are distinct from each other and how models’ ICL performance is mainly driven by the former. ICL can also be viewed as a mixture of meta-learning and retrieval [52]. In that regard, we study the latter perspective to understand its mechanism as a specific instantiation of ICL, with a focus on the retrieval augmentation paradigm.
@@ 3 Background and Preliminaries
An alternative would be to use feature attribution methods [9], as they trace the contribution of each input element to the model’s output prediction. Propagation-based feature attribution [67, 8, 68] especially takes the entire computation path into account, which can be used to characterize attention heads [76] or identify latent concepts [1]. Furthermore, feature attribution is able to estimate causal relations [26] e.g., to automate circuit discovery [18, 70, 22, 30, 32], and thus enables to observe how a specific attention head affects a model’s prediction.
@@ 3.2 Attention Mechanism
While the original formulation of the multi-head attention mechanism [75] concisely summarizes the parallel computation of attention heads, our goal is to isolate their individual contributions. To this end, we reformulate the equations to make the influence of each head more explicit [21, 23]. Let $\mathbf{X}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{S})\in\mathbb{R}^{d\times S}$ denote the matrix of hidden token representations for a sequence of length $S$ with dimension $d$, and suppose our model employs $H$ parallel heads, each of dimension $d_{h}=d/H$. Then, the computation of the multi-head attention layer can be reformulated into $H$ complementary operations, where each head $h$ produces an intermediate attention output $\mathbf{z}_{i}^{h}\in\mathbb{R}^{d_{h}}$:
@@
where $\mathbf{A}^{h}_{i,j}$ is the attention weight of token $i$ attending to token $j$, and $\mathbf{W}_{V}^{h}\in\mathbb{R}^{d_{h}\times d}$ is the per‐head value projection. The final output is obtained by multiplying the intermediate output of each head with their corresponding output projection matrix $\mathbf{W}_{O}^{h}\in\mathbb{R}^{d\times d_{h}}$, followed by summing:
@@ 4 Experimental Setup
Similar to Petroni et al. [58], we consider different types of contextual information to see how they affect in-context and parametric heads. We use oracle contexts as they are always relevant to the question and contain the true answer. In addition, we use counterfactual contexts as they contain information that is usually not seen during pre-training and fine-tuning stages, thus forcing models to rely on the text to answer the question correctly. Oracle context is often not available; therefore we also use Dense Passage Retriever (DPR) [39] with a Wikipedia dump from December 2018 as our retrieval corpus. For simplicity, we only select the top one retrieved document. We show results for oracle and counterfactual contexts in the main paper and retrieved DPR contexts in Appendix B.
@@ 5 Localization of In-Context and Parametric Heads
We examine whether in-context and parametric heads are independent of each other. As expected, ablating parametric heads has little influence to the model’s performance in our open-book setting. Interestingly, this leads to a slight performance increase on NQ-Swap with counterfactual contexts, which suggest that the ablation forces the model to rely more on the given context instead of its own parametric knowledge. Surprisingly, ablating in-context heads in the closed-book setting incurs a non-negligible performance reduction. This is likely due to the influence in-context heads have when processing the input prompt. We explore this in §6.
5.1 Resolving Knowledge Conflicts
We further test whether targeted ablation of parametric and in-context heads also helps to resolve conflicts between contextual and parametric knowledge sources. On NQ-Swap with counterfactual contexts, we selectively ablate heads from each category. Removing in-context heads forces the model to rely more heavily on parametric memory, leading to substantial gains in recall with respect to the original, non-counterfactual answers. Conversely, ablating parametric heads promotes stronger grounding in the provided context as shown before, improving recall for counterfactual open-book answers. We present the full results in Appendix E.
@@ 6 Functional Roles of In-Context and Parametric Heads
[…]
@@ 6.1 Disentangling the Functionality of In-Context Heads
Finally, to obtain the sets of specialized task and retrieval heads, we rank heads by their aggregated relevance and select the top $K$, a hyperparameter determined separately for each experiment.
$\mathcal{H}_{\text{task}}^{K}=\bigl\{\mathrm{argsort}(\rho^{h}_{\text{task}})_{\mathrm{desc}}\bigr\}_{n=1}^{K},\quad\mathcal{H}_{\text{ret}}^{K}=\bigl\{\mathrm{argsort}(\rho^{h}_{\text{ret}})_{\mathrm{desc}}\bigr\}_{n=1}^{K}.$ (9)
[…]
@@ 6.2 Causal Effects of In-Context and Parametric Heads
An important question is whether the heads we identify truly reflect their assigned functionalities. We examine this under controlled conditions and investigate their causal effects on the answer generation.
[…]
@@
Models $\mathcal{H}_{\text{task}}^{40}$ $\mathcal{H}_{\text{param}}^{50}$ $\mathcal{H}_{\text{ret}}^{40}$
@@
6.2.1 How Does This Generalize to Other In-Context Tasks Beyond QA?
To assess whether our method generalizes beyond question answering, we conduct a preliminary investigation on machine translation using the OPUS Europarl dataset [41, 73]. Directly transferring the task heads identified in the QA experiments proved ineffective, indicating that translation relies on a different functional subspace. Nonetheless, by applying our disentanglement procedure of §6.1 to the intensional frame of a translation prompt, we are able to identify a compact set of 15 attention heads that robustly encode translation behavior independent of source or target language. Patching these heads induces zero-shot translation across multiple language pairs, yielding BLEU scores comparable to explicit prompting whereas patching random heads fails to induce any translation behavior. Complete results and experimental details are provided in Appendix F.
@@ 7 Source Tracking
Our experiment in §6.2 demonstrates that retrieval heads reliably perform verbatim copying of text spans when their corresponding attention maps focus on the retrieved tokens. As can be seen on Figure 5, we now aim to investigate if we can (i) detect when retrieval heads initiate the copying process for the first answer token (i.e., whether a token is derived from external contexts rather than from the model parameters), and (ii) accurately localize its position within that context using the attention maps. To this end, we train a linear probe on NQ-Swap with counterfactual contexts. Each retrieval head’s output at the last token’s position $\mathbf{z}_{S}^{h}$ is decoded via logit lens [53], converting each head’s activation into a score for token $t\in\mathbb{N}$ using the model’s unembedding matrix $\mathbf{W}_{U}\in\mathbb{R}^{|V|\times d}$ and layer normalization $\mathrm{LN}(\cdot)$:
@@
Next, we train a probe via linear regression, yielding the weights $\{w_{h}\}_{h\in\mathcal{H}_{\mathrm{ret}}}$. For source localization, we aggregate each head’s attention map using its weight and logit-lens score, and predict the source token index as
$\hat{k}=\arg\max_{j}\sum_{h\in\mathcal{H}_{\mathrm{ret}}}w_{h}\,\mathscr{L}\bigl(\mathbf{z}_{S}^{h}\mid t\bigr)\,A^{h}_{S,j}.$ (11)
For details, please refer to Appendix D. Additionally, we use a standard AttnLRP backward pass from the model output to compute an input heatmap as a baseline for comparison.
@@
Models ROC AUC Localization
Attention AttnLRP
@@ 8 Conclusion
We propose a method to explore the inner workings of ICL for retrieval-augmentation, revealing in-context and parametric heads that operate on the input prompt in a distinct manner and find that in-context heads can specialize into either task or retrieval heads, depending on whether they encode intensional frames or retrieve relevant information. We study the roles of the identified heads by converting them into FVs or modifying their weights, showing how they can affect the generation process. Finally, we present a probe to precisely and efficiently track for knowledge provenance, opening up a path towards more interpretable retrieval-augmented LMs.
Limitations    We focus our investigation on attention heads since they are primarily associated with ICL. However, how they interact with components in MLP modules e.g., knowledge neurons [19] to induce functions remains an open question. Our analyses are also mainly centered on QA (with a minor part on MT). It would be interesting to see if similar mechanisms arise in other tasks. We find that several heads exhibit inhibitory effects, and that some intermediate heads are not exclusively specialized for either in-context or parametric roles — both of which warrant further investigation. In addition, there is a possibility for redundant heads [79], which are yet to be uncovered. We leave this avenue for future work.
@@
Acknowledgements
We thank the anonymous reviewers for their critical reading of our manuscript and their insightful comments and suggestions.
@@ Appendix A Appendix
Llama-3.1-8B-Instruct is released under the Llama 3.1 Community License. Gemma-2-9B-it is released under the Gemma license agreement. Mistral-7B-Instruct-v0.3 is released under Apache 2.0. NQ-Swap and TriviaQA that we use are derived from MRQA [24], which is released under the MIT license. We construct the biography dataset using Wikidata Query Service, [^4] https://query.wikidata.org/ which is available under CC0. The OPUS Europarl dataset is also available under CC0.
QA Datasets
For NQ-Swap, we use the preprocessed data and split available on HuggingFace [^5] https://huggingface.co/datasets/pminervini/NQ-Swap (4,746 examples). The corpus substitution procedure [48] is applied to generate counterfactual contexts. As for TriviaQA, we use the dev split from the MRQA repository [^6] https://github.com/mrqa/MRQA-Shared-Task-2019 (7,785 examples).
@@
ORIG
Vladimir Vapnik was educated in V.A. Trapeznikov Institute of Control Sciences. Vladimir Vapnik was born 06 December 1936. Vladimir Vapnik worked as computer scientist. Vladimir Vapnik was born in Tashkent.
PARA
Vladimir Vapnik was born on 06 December 1936 in Tashkent, a city that would later shape his life’s work. As a young man, Vapnik was fascinated by the potential of machines to learn and adapt. He went on to study at the V.A. Trapeznikov Institute of Control Sciences, where he was exposed to the latest advancements in computer science and artificial intelligence. It was here that Vapnik’s passion for machine learning truly took hold. After completing his studies, Vapnik went on to become a renowned computer scientist, making groundbreaking contributions to the field. His work on support vector machines and the Vapnik-Chervonenkis dimension would have a lasting impact on the development of machine learning algorithms. Throughout his career, Vapnik has been recognized for his innovative thinking and dedication to advancing the field of computer science. His legacy continues to inspire new generations of researchers and scientists.
@@
All the experiments were conducted on 2 x 24 GB RTX4090 and 4 x 40 GB A100. Computing difference score $\mathcal{D}$ to identify in-context and parametric heads takes about 8 hours. Ablating both heads on NQ-Swap and TQA takes about 3 hours for each run. Patching task and parametric FVs takes about 12 hours on average for each model, while the NIAH experiment with retrieval heads takes about 8 hours. Our source tracking experiments consume about 4 hours. The machine translation experiment takes about 6 hours.